OpenStudy (anonymous):

u^1/3 * u^5/3 apparently equals u^2 but I can't figure out how to get to that point.

OpenStudy (oaktree):

When you multiply variables, you add the exponents, right? So we have\[u^{1/3} \times u^{5/3} = u^{1/3 + 5/3} = u^2\]Do you get it now?

OpenStudy (anonymous):

Oh it's so simple! I was converting them into squares before evaluating. Thanks!

OpenStudy (oaktree):

No problem. Thanks for the medal!